Solar constant on earth is 1.36 × 103 W m–2, If Mars is at a distance of 1.52 AU from the
sun, calculate the solar constant on Mars applying Stefans Law.
1
Expert's answer
2018-11-29T12:30:10-0500

The solar constant is a power of solar electromagnetic radiation per unit area. The Stefan’s Law gives

Power=σT^4 A

Here A is an area of Sun surface.

The power of solar electromagnetic radiation per unit area near surface of planet

Solar constant=G_sc=(σT^4 A)/(4πa^2 )

So

(G_sc (Mars))/(G_sc (Earth) )=(a_E^2)/(a_M^2 )=(1 AU)^2/(1.52 AU)^2 =0.433

Finally, the solar constant on Mars

G_sc (Mars)=0.433×G_sc (Earth)=0.433×1.36×〖10〗^3=0.589×〖10〗^3 W∙m^(–2)